# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ 4.4 Wald-type tests for validity of the Michaelis-Menten model
In the context of our MM model (28), we assume $x_{i}>0$, for all $i=1,2,\ldots,n$, as $x_{i}$’s are the substrate concentrations in a reaction. Now, $\beta_{2}=0$ implies that the mean/regression function is simply $\mu(x_{i},\bm{\beta})=\beta_{1}$, a constant, which is independent of the covariate $x_{i}$; Consequently $y_{i}$’s simply become IID observations from a normal distribution with constant mean $\beta_{1}$ and variance $\sigma^{2}$. So, to test the validity of the MM model assumption it is enough to test
@@
Then, for testing $H_{0}$ against the one sided alternative $H_{1}$ in (34), $H_{0}$ will be rejected at level of significance $\gamma$ if the observed value of $\widetilde{W}_{n,\alpha}$ exceeds $z_{\gamma}$, the $(1-\gamma)$-th quantile of $\mathcal{N}(0,1)$. Similarly, one can construct Wald-type tests for testing of the parameter $\beta_{1}$ against one sided or both sided alternatives.
We have conducted simulation studies to study the performance of the Wald-type test for the above discussed testing problem. In particular, we have computed the empirical level (ELevel) and the empirical power (EPower) of the test for various values of the tuning parameter $\alpha$ based on 1000 replications of simulated data; the simulation schemes for computations of ELevel and EPower are taken as follows:
For ELevel: We have simulated samples as per Setup 2, but with $(\beta_{1},\beta_{2},\sigma)=(20,0,1)$ and different sample sizes ($n$). So, $y_{i}$’s are IID $\mathcal{N}(20,1)$, i.e., the MM model is not significant for these (pure) data. For introducing contamination, $e_{c}$ proportion of $y_{i}$’s are randomly selected and each $y_{i}$ is replaced by a random number between $z_{i}$ and $y_{i}$, where $z_{i}=\frac{20x_{i}}{5+x_{i}}$ is the true mean function of an MM model with parameters $\beta_{1}=20$, $\beta_{2}=5$. We have chosen such a contamination scheme, so as to make the (non-robust) estimates of $\beta_{2}$ likely to be larger than 0, so that a non-robust test will reject the null hypothesis more than expected due to the effect of data contamination, although the null hypothesis is true for the actual generated samples. The ideal values of ELevel, measured as the proportion of rejections, should be close to the level of the test, which is taken to be $5\%$ in our experiments.
For EPower: We have again simulated data following Setup 2 but with $(\beta_{1},\beta_{2},\sigma)=(20,1,2)$ and different sample sizes ($n$). For introducing contamination, $e_{c}$ proportion of $y_{i}$’s are randomly selected and each $y_{i}$ is replaced by a random number generated from Uniform$(0,20)$. In this case the null hypothesis is actually false, i.e., the MM model is indeed valid. So, empirical power, measured as the proportion of rejections, should be high (close or equal to 1) for a good stable test.
@@
One can observe from Figure 2 that, under pure data the empirical levels of Wald test (at $\alpha=0$) and all the MDPDE based Wald-type tests with $\alpha>0$ are quite close to each other and fluctuate closely around the desired $5\%$ level of significance. Further, for contaminated data, the empirical levels of the MDPDE based Wald-type tests with moderate or large values of $\alpha$ $(\alpha\geq 0.3)$ tend to stabilize near $5\%$ level of significance, as sample size increases. But the empirical levels of the MLE based Wald test fail to converge to the desired level in the presence of contamination due to its non-robust nature.
@@
Figure 3 illustrates that, under pure data, the empirical powers of the Wald test and MDPDE based the Wald-type tests are all very close to 1 at a moderate sample size of $n=50$ or $60$, and clearly converge to 1 quite fast as the sample size increases. But for contaminated data, the empirical powers of the Wald tests decrease considerably with increase in the proportion of contamination and the rate of consistency slows down significantly. However, the empirical powers of the MDPDE based Wald-type tests with $\alpha\geq 0.3$ decreases only a little at higher contamination proportions but then also they regain their convergence rate to one very quickly as the sample size increases. These figures clearly illustrate significantly improved stability of our proposed MDPDE based Wald-type tests with increasing values of $\alpha>0$ under possible data contamination.
@@
5 On the choice of optimum tuning parameter
@@
It is well studied that the tuning parameter $\alpha$ of the MDPDE acts as a trade-off between robustness and efficiency, in the sense that the robustness of the estimator, under data contamination, increases at the cost of loss in (asymptotic) efficiency under pure data with increasing values of $\alpha$. The same is also clearly observed for the MDPDEs and the associated Wald-types tests under our NLR model (3) through the extensive numerical illustrations provided in the previous sections. So, while applying these DPD based estimation and testing methods to a real dataset, it is very important to choose an optimal value of $\alpha$ providing the best possible ‘compromise’ between robustness and efficiency depending on the (unknown) amount of contamination in the given dataset.
@@
Under the setup of IID data, Warwick and Jones [39] suggested to use the optimal value of $\alpha$ that minimizes an asymptotic approximation of the MSE of the MDPDEs. Ghosh and Basu [15] extended this idea for the general INH setup, with detailed illustrations for the linear regression model. Following the same idea, an optimal value of $\alpha$ for our NLR model (3) can be obtained by minimizing an estimate of the quantity
$E(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta}^{*})^{T}(\bm{\widehat{\theta}_{\alpha}}-\bm{\theta}^{*})=(\bm{\theta_{\alpha}}-\bm{\theta}^{*})^{T}(\bm{\theta_{\alpha}}-\bm{\theta}^{*})+n^{-1}Trace(\bm{\Psi_{n}^{-1}(\bm{\theta_{\alpha}})\Omega_{n}(\bm{\theta_{\alpha}})\Psi_{n}^{-1}(\bm{\theta_{\alpha}})}),$
@@
where $\bm{\theta}^{*}$ is the target (true) parameter value, $\bm{\theta_{\alpha}}$ is the minimum DPD functional given in (16) and $\bm{\Psi_{n},\Omega_{n}}$ are as given in (11). We may estimate this MSE by replacing $\bm{\theta_{\alpha}}$ by $\bm{\widehat{\theta}_{\alpha}}$ and $\bm{\theta}^{*}$ by a pilot estimator, and then minimize the resulting estimated MSE by grid search over a suitably chosen grid of $\alpha\in[0,1]$. This method crucially depends on the choice of the pilot estimator, although it is shown to work remarkably well for a properly chosen pilot.
@@
In order to remove the dependence on the pilot estimate, Basak et al. [4] proposed an iterative version of the above procedure, namely the Iterated Warwick-Jones (IWJ) method, where they have proposed to use the above procedure iteratively with the pilot estimate being taken as the MDPDE with the tuning parameter obtained in the previous iteration (as optimal). They have numerically illustrated that, starting with the MDPDE with any $\alpha>0$ as the pilot estimate, their IWJ algorithm eventually converges (very fast) to an optimal value of $\alpha$, which is independent of the initial choice of the pilot estimate. We have followed this IWJ approach in order to find the optimal values of $\alpha$ in all our real data examples presented in the following section.
@@ 6 Real data examples
Here we present the application of our proposed robust inference methodologies to fit the MM model (28) to some interesting real-life datasets, and compare them with the fits obtained by the existing robust estimators (except for the MOM estimator for the reason discussed in Section 4.3). However, for real data, we are unaware of the exact amount of contaminated observations and hence the prediction accuracy of a fitted model is assessed by the $p\%$ trimmed average prediction error (TAPE), obtained after the largest $p\%$ prediction errors (squared residuals) are being trimmed.
@@ 6.1 ONPG data
Our first example is the OPNG dataset developed by Aledo [1]. A batch of second year biochemistry students of the University of Málaga, Spain, using $\beta$-galatosidase (EC. 3.2.1.23) as an enzyme model, carried out different experiments on the hydrolysis of o-nitrophenyl-$\beta$-D-galactopyranoside (ONPG) to galactose and o-nitrophenol, which is an enzyme-catalyzed reaction. Our data, obtained from the R package renz [1], consist of observations from 10 such reactions with different ONPG concentrations where the initial rates (velocity) of these 10 reactions were recorded separately by 8 groups of students. Here, for our illustration purpose, we consider the data obtained by the second group of students, which contains some outlying observations (among the total of $n=10$ observations).
Since these observations are obtained from enzyme-catalyzed reactions, we have fitted the MM model (28) to these data with the initial reaction rates (velocities) as the response and the ONPG concentration as the covariate, using our proposed MDPDEs as well as its robust competitors. The parameter estimates and their (estimated) asymptotic standard errors are given in Table 6 along with the $30\%$ TAPE based on the fitted models for each method. The resulting fitted lines for some selected methods are also shown in Figure 4, along with the scatter plot of the actual observations. For the selection of optimum $\alpha$, we have used the IWJ method as described in Section 5, which yields $\widehat{\alpha}=0.26$, in just one or two iterations for any choice of pilot $\alpha>0$.
Table 6: Parameter estimates of the MM model fitted to the ONPG data, along with their asymptotic standard errors (in parenthesis) and 30% TAPE. [For MDPDE, $\widehat{\alpha}=0.26$ is the optimum choice obtained by IWJ method]
@@
We can see from Table 6 and Figure 4 that the classical OLS fit is heavily affected by the presence of outliers in the data. The KPS method is ignoring the rightmost 2 observations even though they are not outlying observations. Our proposed MDPDE with the optimum value of $\alpha=0.26$ provides much better and stable parameter estimates with lower prediction error for most of the data; in fact, all the values of $\alpha\geq\widehat{\alpha}$ yield similar estimators and model fits. MDPDE with $\alpha=0.1$ and Tukey’s M-estimator also yield reasonable model fits but higher TAPE compared to the MDPDE with $\alpha=0.26$.
@@
Subsequently, we have also tested for the validity of the Michaelis-Menten model (i.e, test of $\beta_{2}$ following the hypothesis in (34)) for this dataset using the classical Wald test and the proposed MDPDE based Wald-type test. The observed values of the test statistic $\widetilde{W}_{n,\alpha}$, described in (35) in Section 4.4, are given in Table 7 along with the associated (one-sided) p-values.
@@
It is evident that the classical Wald test (at $\alpha=0$) fails to reject the null hypothesis in (34) at $5\%$ level of significance and hence fails to justify the validity of the MM model for this particular dataset, due to the presence of outliers. However, our MDPDE based Wald-type tests with $\alpha\geq 0.1$ provide robust inferences by rejecting the null hypothesis in (34) and hence justifying the validity of the MM model. In fact, these Wald-type tests at the optimum value of $\alpha$ and beyond (i.e., for $\alpha\geq 0.26$) all reject the null hypothesis at any reasonable level of significance, successfully suppressing any effect of outlying observations present in the dataset.
@@ 6.2 Drug concentration data
Our second dataset, taken from Tabatabai et al. [36], contains a set of responses for 7 different concentrations of an agonist in a functional assay. Among the 7 observations, the 5-th one is clearly seen to be an outlier; see the scatter plot in Figure 5. Once again we have fitted the MM model (28) for this dataset using the proposed MDPDE, as well as its existing competitors. The parameter estimates are given in the Table 8 along with the asymptotic standard errors and $30\%$ TAPEs for the fitted models, while the model fits are illustrated in Figure 5. The optimum value of $\alpha$ in our MDPDEs, using the IWJ method, turns out to be $\widehat{\alpha}=0.39$ for this dataset.
Like the previous example, here also we can see that the proposed MDPDE at the optimum $\alpha=0.39$ (in fact for all $\alpha\geq 0.39$) yields robust parameter estimates and model fits ignoring the effects of the outlier. Tukey’s M-estimator still remains a close contender with similar parameter estimate with a lower TAPE, but form the Figure 5 it can be clearly observed that it does not provide a fit as good as the MDPDEs with $\alpha\geq 0.39$. Also, the KPS estimators for both $w=0.05$ and $w=1.5$ are heavily affected by the presence of outlier as are the OLS and the MDPDE with $\alpha=0.1$.
We have also performed the Wald-type tests for testing the validity of the MM model for this dataset. Interestingly, all these tests (for all $\alpha\geq 0$) reject the null hypothesis of insignificance of the MM model with a p-value $\leq 0.001$ in all cases. The outlier has no effect on the validity of MM model for this example, and hence all values of $\alpha$ lead to the same inference as the classical Wald test.
Table 8: Parameter estimates of the MM model fitted to the Drug concentration data, along with their asymptotic standard errors (in parenthesis) and 30% TAPE. [For MDPDE, $\widehat{\alpha}=0.39$ is the optimum choice obtained by IWJ method]

### v2

@@ 4.4 Wald-type tests for validity of the Michaelis-Menten model
In the context of our MM model (28), we assume $x_{i}>0$, for all $i=1,2,\ldots,n$, as $x_{i}$’s are the substrate concentrations in a reaction. Now, $\beta_{2}=0$ implies that the mean/regression function is simply $\mu(x_{i},\bm{\beta})=\beta_{1}$, a constant, which is independent of the covariate $x_{i}$; Consequently, $y_{i}$’s simply become IID observations from a normal distribution with constant mean $\beta_{1}$ and variance $\sigma^{2}$. So, to test the validity of the MM model assumption, it is enough to test
@@
Then, for testing $H_{0}$ against the one-sided alternative $H_{1}$ in (34), $H_{0}$ will be rejected at level of significance $\gamma$ if the observed value of $\widetilde{W}_{n,\alpha}$ exceeds $z_{\gamma}$, the $(1-\gamma)$-th quantile of $\mathcal{N}(0,1)$. Similarly, one can construct Wald-type tests for testing of the parameter $\beta_{1}$ against one-sided or both-sided alternatives.
We have conducted simulation studies to study the performance of the Wald-type test for the above-discussed testing problem. In particular, we have computed the empirical level (ELevel) and the empirical power (EPower) of the test for various values of the tuning parameter $\alpha$ based on 1000 replications of simulated data; the simulation schemes for computations of ELevel and EPower are taken as follows:
For ELevel: We have simulated samples as per Setup 2, but with $(\beta_{1},\beta_{2},\sigma)=(20,0,1)$ and different sample sizes ($n$). So, $y_{i}$’s are IID $\mathcal{N}(20,1)$, i.e., the MM model is not significant for these (pure) data. To introduce contamination, $e_{c}$ proportion of $y_{i}$’s are randomly selected and each $y_{i}$ is replaced by a random number between $z_{i}$ and $y_{i}$, where $z_{i}=\frac{20x_{i}}{5+x_{i}}$ is the true mean function of an MM model with parameters $\beta_{1}=20$, $\beta_{2}=5$. We have chosen such a contamination scheme, to make the (non-robust) estimates of $\beta_{2}$ likely to be larger than 0, so that a non-robust test will reject the null hypothesis more than the expected due to the effect of data contamination, although the null hypothesis is true for the actual generated samples. The ideal values of ELevel, measured as the proportion of rejections, should be close to the level of the test, which is taken to be $5\%$ in our experiments.
For EPower: We have again simulated data following Setup 2 but with $(\beta_{1},\beta_{2},\sigma)=(20,1,2)$ and different sample sizes ($n$). To introduce contamination, $e_{c}$ proportion of $y_{i}$’s are randomly selected and each $y_{i}$ is replaced by a random number generated from Uniform(0,20)(0,20). In this case, the null hypothesis is actually false, i.e., the MM model is indeed valid. So, empirical power, measured as the proportion of rejections, should be high (close or equal to 1) for a good stable test.
@@
Figure 2 shows that, under pure data, the empirical levels of the Wald test (at $\alpha=0$) and all the MDPDE-based Wald-type tests with $\alpha>0$ are quite close to each other and fluctuate closely around the desired $5\%$ level of significance. Further, for contaminated data, the empirical levels of the MDPDE-based Wald-type tests with moderate or large values of $\alpha$ $(\alpha\geq 0.3)$ tend to stabilize near $5\%$ level of significance, as sample size increases. But the empirical levels of the MLE-based Wald test fail to converge to the desired level under contamination due to its non-robust nature.
Figure 3 illustrates that, under pure data, the empirical powers of the Wald test and MDPDE-based Wald-type tests are all very close to 1 at a moderate sample size of $n=50$ or $60$, and clearly converge to 1 quite fast as the sample size increases. But for contaminated data, the empirical powers of the Wald tests decrease considerably with increase in the proportion of contamination, and the rate of consistency slows down significantly. However, the empirical powers of the Wald-type tests with $\alpha\geq 0.3$ decrease only a little at higher contamination proportion, but then they also regain their convergence rate to one very quickly as the sample size increases. These figures clearly illustrate significantly improved stability of our proposed MDPDE-based Wald-type tests with increasing values of $\alpha>0$ under possible data contamination.
5 On the choice of optimum tuning parameter $\alpha$
@@
It is well studied that the tuning parameter $\alpha$ of the MDPDE acts as a trade-off between robustness and efficiency — the robustness of the estimator under data contamination increases at a minimal cost of the loss in (asymptotic) efficiency under pure data as $\alpha$ increases. The same is also observed for the MDPDEs and the associated Wald-type tests under our NLR model (3) across all the simulation results discussed in the previous sections. So, while applying these DPD-based estimation and testing methods to a real dataset, it is very important to choose an optimal value of $\alpha$ providing the best possible ‘compromise’ between robustness and efficiency, depending on the (unknown) amount of contamination in the given dataset.
@@
Under the setup of IID data, Warwick and Jones [42] suggested using the optimal value of $\alpha$ that minimizes an asymptotic approximation of the MSE of the MDPDEs. Ghosh and Basu [14] extended this idea for the general INH setup, with detailed illustrations for the linear regression model. Following the same idea, an optimal value of $\alpha$ for our NLR model can be obtained by minimizing an estimate of the quantity
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E(\bm{\widehat{\theta}}_{\alpha}-\bm{\theta}^{*})^{T}(\bm{\widehat{\theta}}_{\alpha}-\bm{\theta}^{*})=(\bm{\theta}_{\alpha}-\bm{\theta}^{*})^{T}(\bm{\theta}_{\alpha}-\bm{\theta}^{*})+n^{-1}Trace\left(\bm{\Psi}_{n}^{-1}(\bm{\theta}_{\alpha})\bm{\Omega}_{n}(\bm{\theta}_{\alpha})\bm{\Psi}_{n}^{-1}(\bm{\theta}_{\alpha})\right),$
@@
where $\bm{\theta}^{*}$ is the target (true) parameter value, $\bm{\theta}_{\alpha}=\bm{T}_{\alpha}(\bm{G})$ is the minimum DPD functional given in Eq. (15) and $\bm{\Psi_{n},\Omega_{n}}$ are as given in Eq. (10). We may estimate this MSE by replacing $\bm{\theta}_{\alpha}$ by $\bm{\widehat{\theta}}_{\alpha}$ (as defined in Section 2.1) and $\bm{\theta}^{*}$ by a pilot estimator, which can then be minimized by the grid search over a suitably chosen grid of $\alpha\in[0,1]$. However, this method crucially depends on the choice of the pilot estimator $\bm{\theta}^{*}$, although it is shown to work very well for a properly chosen pilot.
@@
To remove the dependence on the pilot estimate, Basak et al. [4] proposed an iterative version of the above procedure, namely the Iterated Warwick-Jones (IWJ) method, where one may use the above procedure iteratively with the pilot estimate being taken as the MDPDE with the optimum tuning parameter value obtained in the previous iteration. They have numerically illustrated that, starting with the MDPDE with any $\alpha>0$ as the pilot estimate in the first iteration, their IWJ algorithm eventually converges (very fast) to an optimal value of $\alpha$, which is independent of the initial choice of the pilot estimate. We have followed this IWJ approach to find the optimal values of $\alpha$ in all our real data examples presented in the following section.
@@ 6 Real data examples
Here we present the application of our proposed robust inference methodologies to fit the MM model (28) to some interesting real-life datasets, and compare them with the fits obtained by the existing robust estimators (except for the MOM estimator for the reason discussed in Section 4.3). In this section, two examples are presented having strong outlier effect, and two additional examples with mild or no outliers are discussed in Web Appendix D. For computing the MDPDEs in all these data examples, we have again minimized the objective function in (5) using optim in R (with the BFGS method); several initial values around the LSE are tried, and the estimate leading to the minimum value of the objective function is finalized and reported.
Further, since we are unaware of the exact amount of contaminated observations for these datasets, the prediction accuracy of the fitted model is assessed by the $p\%$ trimmed average prediction error (TAPE), obtained after the largest $p\%$ prediction errors (squared residuals) are being trimmed. We have reported the results with 30% trimming as a conservative choice of $p$ for the two datasets with strong outlier effects, and 20% trimming for the others having milder contamination.
@@ 6.1 ONPG data
Our first example is the ONPG dataset, developed by Aledo [1]. A batch of second-year biochemistry students of the University of Málaga, Spain, using $\beta$-galactosidase (EC. 3.2.1.23) as an enzyme model, carried out different experiments on the hydrolysis of o-nitrophenyl-$\beta$-D-galactopyranoside (ONPG) to galactose and o-nitrophenol, which is an enzyme-catalyzed reaction. Our data, obtained from the R package renz [1], consist of observations from 10 such reactions with different ONPG concentrations, where the initial rates (velocity) of these 10 reactions were recorded separately by 8 groups of students. Here, for illustrative purposes, we consider the data obtained by the second group of students, which contains some outlying observations (among the total of $n=10$ observations).
Since these observations are obtained from enzyme-catalyzed reactions, we fitted the MM model (28) to these data with the initial reaction rates (velocities) as the response and the ONPG concentration as the covariate, using our proposed MDPDEs as well as its robust competitors. The parameter estimates and their (estimated) asymptotic standard errors are given in Table 3 along with the $30\%$ TAPE based on the fitted models for each method. The resulting fitted lines for some selected methods are also shown in Figure 4, along with the scatter plot of the actual observations. For the selection of optimum $\alpha$, we used the IWJ method as described in Section 5, which yields $\widehat{\alpha}=0.26$, in just one or two iterations for any choice of pilot $\alpha>0$.
Table 3: Parameter estimates of the MM model fitted to the ONPG data, along with their asymptotic standard errors (in parentheses) and 30% TAPE. [For MDPDE, $\widehat{\alpha}=0.26$ is the optimum choice obtained by IWJ method]
@@
We can see from Table 3 and Figure 4 that the classical OLS fit is heavily affected by the presence of outliers in the data. The KPS method is ignoring the rightmost 2 observations even though they are not outlying observations. Our proposed MDPDE with the optimum value of $\alpha=0.26$ provides much better and stable parameter estimates with lower prediction error for most of the data; in fact, all the values of $\alpha\geq\widehat{\alpha}$ yield similar estimators and model fits. MDPDE with $\alpha=0.1$ and Tukey’s M-estimator also yield reasonable model fits but higher TAPE compared to the MDPDE with $\alpha=0.26$ (optimum choice).
The robust MDPDEs (at the optimum $\alpha=0.26$) of the MM model parameters suggest a significantly lower value of $V_{max}$ (128.515 vs 165.004) and $K_{m}$ (2.582 vs 12.350) compared to the classical OLS estimates, indicating an upward bias in the classical non-robust fit. This is potentially caused by a clear outlier present in the data having a significantly low recorded initial reaction velocity. This smaller estimated $K_{m}$ indicates a higher substrate affinity, suggesting the $\beta$-galactosidase reaches the half-maximal velocity at a lower ONPG concentration than the value predicted by the OLS estimate. As a result, our estimate enhances the reliability in the enzyme-kinetic modelling of $\beta$-galactosidase enzyme with ONPG substrate, avoiding overestimation of the enzyme efficiency that might otherwise lead to inaccurate decisions during industrial bio-catalysis scale-up.
Subsequently, we also tested for the validity of the Michaelis-Menten model (i.e, test of $\beta_{2}$ following the hypothesis in (34)) for this dataset using the classical Wald test and the proposed MDPDE-based Wald-type test. The observed values of the test statistic $\widetilde{W}_{n,\alpha}$, described in (35), are given in Table 4 along with the associated (one-sided) p-values. It is evident that the classical Wald test (at $\alpha=0$) fails to reject the null hypothesis in (34) at $5\%$ level of significance and hence fails to justify the validity of the MM model for this particular dataset, due to the presence of outliers. However, our MDPDE-based Wald-type tests with $\alpha\geq 0.1$ provide robust inferences by rejecting the null hypothesis in (34) and hence justifying the validity of the MM model. In fact, these Wald-type tests at the optimum $\alpha$ and beyond (i.e., for $\alpha\geq 0.26$) all reject the null hypothesis at any reasonable level of significance, successfully suppressing any effect of outlying observations.
@@ 6.2 Drug concentration data
Our second dataset, taken from Tabatabai et al. [39], contains a set of responses for 7 different concentrations of an agonist in a functional assay. Among the 7 observations, the 5th one is clearly seen to be an outlier; see the scatter plot in Figure 5. Once again, we have fitted the MM model (28) for this dataset using the proposed MDPDE, as well as its existing competitors. The parameter estimates are given in the Table 5 along with the asymptotic standard errors and $30\%$ TAPEs for the fitted models, while the model fits are illustrated in Figure 5. The optimum value of $\alpha$ in our MDPDEs, using the IWJ method, turns out to be $\widehat{\alpha}=0.39$ for this dataset.
Table 5: Parameter estimates of the MM model fitted to the Drug concentration data, along with their asymptotic standard errors (in parentheses) and 30% TAPE. [For MDPDE, $\widehat{\alpha}=0.39$ is the optimum choice obtained by IWJ method]